OpenStudy (anonymous):

Please help!!! find the equation of the line containing the given points (3,4) and (9,9)

OpenStudy (ash2326):

@dawn354215 equation of a line passing through two points (x1, y1) and (x2, y2) is given as \[\frac{y-y2}{x-x2}=\frac{y2-y1}{x2-x1}\] Here (x1, y1)= (3,4) and (x2, y2) =(9, 9) Can you try now?

OpenStudy (anonymous):

Find the gradient first using change in Y/change in X. As explained above then use one of the given points to find the equation. \[y-4\div x-3\] = Gradient and cross multiply